A laboratory technician has made up a stock mixture of two chemicals, X and Y, ready for use by all of the groups in Year 9 at a High School. He has mixed 6.0 kg which is 1/4 of X and 3/4 of Y. He realises that he has used the wrong mix and, rather than waste the chemicals he has used so far, he is going to add some more of chemical X so that the mix consists of 40% of X and 60% of Y.
What mass of chemical X must he add?
1
Expert's answer
2019-07-22T03:25:18-0400

There are 6*0.25=1.5 kg of chemical X and 6*0.75=4.5 kg of chemical Y in the initial mix.

Let x be a mass of chemical X which must be added.

Thus, "\frac{1.5+x}{4.5}=\frac{40}{60}=\frac{2}{3}."

So, "4.5+3x=9" or x=1.5.

He must add1.5 kg of chemical X.